# Dust evolution during the protostellar collapse: influence on the coupling between the neutral gas and the magnetic field

### v1

@@ Abstract
Context. The degree of coupling between the magnetic field and the gas during the collapsing phase of star-forming cores is strongly affected by the dust size distribution, which is expected to evolve by means of coagulation, fragmentation and other collision outcomes.
Aims. We aim to investigate the influence of key parameters on the evolution of the dust distribution as well as on the magnetic resistivities during the protostellar collapse.
Methods. We perform collapsing single zone simulations with shark. The code computes the evolution of the dust distribution, accounting for different grain growth and destruction processes, the grain collisions being driven by brownian motion, turbulence and ambipolar drift. It also computes the charges carried by each grain species, the ion and electron densities as well as the magnetic resistivities.
Results. We find that the dust distribution significantly evolves during the protostellar collapse, shaping the magnetic resistivities. The peak size of the distribution, the population of small grains and consequently the magnetic resistivities are controlled by both coagulation and fragmentation rates. Under standard assumptions, the small grains coagulate very early as they collide by ambipolar drift, yielding magnetic resistivities orders of magnitude away from the non-evolving dust case. In particular, the ambipolar resistivity $\eta_{\mathrm{AD}}$ is very high prior to $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$, and as a consequence magnetic braking should be ineffective. In this case, large size protoplanetary disks should result, which is inconsistent with recent observations. To alleviate this tension, we identify mechanisms to reduce the ambipolar resistivity during the protostellar collapse. The most promising are namely: electrostatic repulsion and grain-grain erosion.
Conclusions. The evolution of the magnetic resistivities during the protostellar collapse and consequently the shape of the magnetic field in the early life of the protoplanetary disk strongly depends on the possibility to repopulate the small grains or to prevent their early coagulation. Therefore, it is crucial to better constrain the collision outcomes and the dust grain elastic properties, especially the grain surface energy from both theoretical and experimental endeavors.
@@
Hydrodynamics; Magnetohydrodynamics (MHD); Turbulence; Protoplanetary disks; Planets and satellites formation.
@@ 1 Introduction
Interstellar dust plays a significant role in many astrophysical environments, in particular, in the context of star formation. Although representing on average $1\%$ of the diffuse ISM (interstellar medium) mass, the wide spectrum of dust grain sizes affects the optical properties of the medium, and consequently the heating/cooling processes at work (McKee & Ostriker, 2007). Besides, dust grains are at the heart of many molecule formation processes (including $\mathrm{H_{2}}$ Gould & Salpeter, 1963). In addition, they are the fundamental bricks of planets, which are believed to form rapidly in less than 1 Myr (Testi et al., 2014). The pathway to planetesimals and thereafter planets remains poorly understood. Better describing the dust evolution during the protostellar collapse would allow one to provide more realistic dust initial conditions in protoplanetary disks. Finally, dust grains affect the ionization degree of the medium and can become the main charge carriers during the protostellar collapse as they allow electron/ion recombination and electron capture at their surface (Marchand et al., 2019). As such, they control the degree of coupling between the gas and the magnetic field (Nakano et al., 2002), via the ambipolar, Hall and ohmic magnetic resistivities and consequently the evolution of the angular momentum through magnetic braking (Mouschovias, 1991). Indeed, dust plays a crucial role in shaping the magnetic field during the protostellar collapse and thus in influencing the structure of the resulting protoplanetary disks (Wurster et al., 2016, 2018; Hennebelle et al., 2020).
@@
The dust size spectrum is observed, or at least modelled as a power law MRN (Mathis-Rumple-Nordsiek) distribution in the diffuse ISM (Mathis et al., 1977). It was noticed that modifying the population of very small grains in the protostellar collapse simulations radically affects the magnetic resistivities and consequently the formation of protoplanetary disks (Zhao et al., 2016; Marchand et al., 2016). In addition, the dust size distribution is expected to evolve as dust grains undergo coagulation and fragmentation processes (Dominik & Tielens, 1997; Ormel et al., 2009). Following such an evolution of the dust distribution is very challenging but essential to correctly describe the non-ideal MHD effects since the coupling between the grains, the gas and the magnetic field depends directly on the total dust cross section. Indeed Guillet et al. (2020), Silsbee et al. (2020) and Lebreuilly et al. (2023) showed that a coagulating dust distribution would affect the resistivities during the protostellar collapse. So far, only few studies have accounted for dust coagulation in 2D (Vorobyov & Elbakyan, 2019; Tu et al., 2022) or 3D (Tsukamoto et al., 2021; Bate, 2022; Tsukamoto et al., 2023). Recently Marchand et al. (2023) have performed, for the first time, 3D simulations of the protostellar collapse and the early evolution of the protoplanetary disk while including and computing self-consistently the magnetic field from the magnetic resistivities. The dust coagulation (fragmentation was not included) is modelled at low computational cost with a turbulent kernel (as described in Marchand et al., 2021), and exerts an indirect feedback on the gas through the magnetic field, allowing to better describe the effect of dust growth on non-ideal MHD effects. Similarly, Tsukamoto et al. (2023) used a dust single-size approximation and describe the co-evolution of dust grains and protoplanetary disks. Fragmentation has been introduced by Lebreuilly et al. (2023) and Kawasaki et al. (2022), and some parameters (turbulent intensity and velocity threshold for fragmentation) have been investigated in more detail in Kawasaki & Machida (2023). They show that the interplay between coagulation and fragmentation plays a critical role in regard to the evolution of the magnetic resistivities. Indeed, those studies showed that fragmentation could potentially repopulate the small grains at intermediate and high densities, consequently allowing the Ohm and Hall resistivities to rise. The ambipolar resistivity $\eta_{\mathrm{AD}}$ is reduced to some extent prior to $n_{\mathrm{H}}<10^{10}\mathrm{cm^{-3}}$ and enhanced thereafter. Besides, every study accounting for dust evolution produces magnetic resistivities which remain remarkably away from the ones obtained with a non-evolving MRN distribution. This is even more striking when including ambipolar diffusion as a source of grain-grain collision, which is very efficient at depleting the smallest grains. For instance, as a consequence of small grain depletion by coagulation, the ambipolar resistivity $\eta_{\mathrm{AD}}$ exhibits very high values at low and intermediate densities. This would have a profound impact on the properties of the resulting protoplanetary disks. Given that recent observational surveys including CALYPSO (Maury et al., 2019) showed that large disks ($>60\ \mathrm{AU}$) are rare, serious doubts are cast on the likeliness of such high ambipolar resistivity values.
@@
To shed light on the impact of dust evolution on the magnetic resistivities during the protostellar collapse, we conducted a thorough investigation of the influence of key parameters and physical effects through single-zone collapse numerical simulations. In particular, we investigate the role of dust grain properties (grain surface energy, intrinsic density, monomer size), collision outcomes (fragments distribution, maximum fragment size, grain-grain erosion, sticking efficiency, electrostatic repulsion) and the collapse properties (dust-to-gas ratio, turbulence intensity, ambipolar drift intensity, initial magnetic field strength and ionization rate induced by cosmic-rays). However, we focus only on the most influential of the aforementioned parameters and strive to identify mechanisms allowing to mitigate the sharp rise in ambipolar resistivity $\eta_{\mathrm{AD}}$.
@@
In Sect. 2 and 3, we present the dust model and the numerical methods. Then, in Sect. 4 and 5, we share some of our results which are discussed in Sect. 6. Finally, Sect. 7 provides concluding remarks, and additional material can be found in the Appendix sections.
@@ 2.1 Sources of dust relative velocities
We present here the different sources of relative velocities which drive dust evolution. In the rest of the paper, the subscripts $i$ and $j$ will refer to the two colliding grains.
@@ 2.1.1 Brownian motion
Dust grains are prone to Brownian motion, which is a thermal process that induces relative velocities as per
@@ 2.1.2 Ambipolar diffusion
where $t_{\mathrm{s,i}}$ is the stopping time, representing the coupling of a grain to the gas (Epstein, 1924)
@@
where $\rho$ and $c_{\mathrm{s}}\equiv\sqrt{\gamma P/\rho}$ are respectively the mass density of the gas and the sound speed. The number of charges carried by a grain increases slower with respect to grain size than the grain mass does. Therefore, larger grains are associated to lower gyromagnetic frequencies, i.e are poorly coupled to the magnetic field. Besides, they are weakly coupled to the gas as well, as suggested by the size dependency of the stopping time. Nonetheless, the stopping time rises less rapidly with respect to the grain size than the inverse of the gyromagnetic frequency (the gyromagnetic timescale) does, leading in fine to lower Hall factors for larger grains.
As in Spitzer (1941), we define the ambipolar diffusion velocity as
@@
where $c$ is the speed of light, $\eta_{\mathrm{AD}}$ is the ambipolar resistivity (in unit $\mathrm{s}$). The curl of the magnetic field is approximated as $\frac{|\mathbf{B}|}{\lambda_{\mathrm{J}}}$ where $\lambda_{\mathrm{J}}$ is the Jeans length. Following Guillet et al. (2007, 2020), we thus have
@@
where $\delta_{\mathrm{AD}}$ is the amibipolar drift intensity, taken equal to unity if not specified otherwise. We note that the larger the ambipolar resistivity $\eta_{\mathrm{AD}}$, the stronger the ambipolar drift velocity. Finally, the ambipolar drift velocity between two dust grains is given by
@@
Ambipolar drift allows the largest grains to collect the smallest ones. Therefore, ambipolar diffusion is a process very effective at removing the smallest grains, but not at increasing significantly the maximum size of the distribution (Lebreuilly et al., 2023). Importantly, we stress that the ambipolar resistivity $\eta_{\mathrm{AD}}$ and the ambipolar drift intensity $\delta_{\mathrm{AD}}$ are two different quantities, which should not be confused in the rest of the paper.
@@ 2.1.3 Turbulence
The Stokes number of a grain is defined as $\mathrm{St_{i}}=t_{\mathrm{s,i}}/t_{\mathrm{ff}}$, where the stopping time $t_{\mathrm{s,i}}$ of the largest of the two grains determines the three coupling regimes defined in the framework of Ormel et al. (2009). The differential turbulent induced velocity writes
@@
where $V_{\mathrm{g}}=\sqrt{\theta}c_{\rm{s}}$ and $\beta_{\mathrm{i,j}}=3.2-(1+x_{i,j})+\frac{2}{1+x_{i,j}}(\frac{1}{2.6}+\frac{x_{i,j}^{3}}{1.6+x_{i,j}})$, $x_{j,i}$ being the ratio between $\mathrm{St}_{j}$ and $\mathrm{St}_{i}$. The parameter $\theta$ is the turbulence intensity. The usual value taken in the literature in the context of a gravitational collapse is $\theta=\frac{3}{2}$ (Guillet et al., 2020).
When the stopping time of the largest grain is comprised between the dissipation timescale and the injection timescale (i.e., turnover time of the largest eddies), the contribution from class II eddies acts as a random kick in the grain motion, allowing to efficiently collide grains of similar size (see Ormel & Cuzzi, 2007, for a detailed definition of the class I, II and III eddies). This is the intermediate regime, which in the context of the protostellar collapse, is the dominant regime (Marchand et al., 2021).
Note that the turbulent source of relative velocities is commonly found to be the most efficient at growing large grains and significantly increasing the maximum size of the dust distribution (Silsbee et al., 2020; Guillet et al., 2020; Lebreuilly et al., 2023).
@@ 2.2 Dust evolution
The different sources of relative velocity presented here lead to grain collisions that can result in various outcomes. In this work, we only account for grain coagulation (perfect and imperfect sticking) and fragmentation, and we also investigate the influence of grain-grain erosion with a simplistic approach. Many other mechanisms have been identified with laboratory experiments, such as gas-grain erosion, abrasion, compaction and bouncing, mass transfer and rotational disruption (see Wurm & Teiser, 2021, for more details). The grains are considered perfectly spherical, with a fixed and unique intrinsic density.
@@ 2.2.1 Coagulation
The equation of mass conservation for the grains writes
@@
The subscript $k$ designates one specific grain among $\mathcal{N}$ species. Here, the term species relates to a grain of given size. The source term $S_{k,\mathrm{growth}}$ accounts for coagulation and fragmentation events and is computed according to the Smoluchowski equation (Smoluchowski, 1916)
@@
where $\Delta v_{i,j}$ is the differential velocity between grains $i$ and $j$, of size $s_{i}$ and $s_{j}$, defined as the quadratic sum of the sources of differential velocity detailed in Sect. 2.1. The $\sqrt{\frac{8}{3\pi}}$ pre-factor comes from considering that grains relative velocities along the three x, y and z axis are Gaussian variables of variance $\Delta v_{i,j}^{2}/3$ (Guillet et al., 2020; Marchand et al., 2021). $E_{\mathrm{kinetic}}=\frac{1}{2}\mu_{i,j}\Delta v_{i,j}^{2}$ is the kinetic energy of the colliding grains. $\mu_{i,j}\equiv m_{i}m_{j}/\left(m_{i}+m_{j}\right)$ is the reduced mass of the two grains. $E_{\mathrm{Coulomb}}=q_{\mathrm{i}}q_{\mathrm{j}}/4\pi\epsilon_{\mathrm{vacuum}}\left(s_{\mathrm{i}}+s_{\mathrm{j}}\right)$ is the Coulomb electrostatic energy and $\epsilon_{\mathrm{vacuum}}$ the vacuum dielectric permittivity. The first term in the Kernel accounts for the cross section modification induced by the electrostatic interaction between the two colliding grains (Akimkin et al., 2023). It is included only in the simulations for which electrostatic repulsion is explicitly investigated.

### v2

# Dust evolution during a protostellar collapse: Influence on the coupling between the neutral gas and magnetic field

@@ Abstract
Context. The extent of the coupling between the magnetic field and the gas during the collapsing phase of star-forming cores is strongly affected by the dust size distribution, which is expected to evolve by means of coagulation, fragmentation, and other collision outcomes.
Aims. We aim to investigate the influence of key parameters on the evolution of the dust distribution, as well as on the magnetic resistivities during protostellar collapse.
Methods. We performed a set of collapsing single-zone simulations with shark. The code computes the evolution of the dust distribution, accounting for different grain growth and destruction processes, with the grain collisions being driven by brownian motion, turbulence, and ambipolar drift. It also computes the charges carried by each grain species and the ion and electron densities, as well as the magnetic resistivities.
Results. We find that the dust distribution significantly evolves during the protostellar collapse, shaping the magnetic resistivities. The peak size of the distribution, population of small grains, and, consequently, the magnetic resistivities are controlled by both coagulation and fragmentation rates. Under standard assumptions, the small grains coagulate very early as they collide by ambipolar drift, yielding magnetic resistivities that are many orders of magnitude apart from the non-evolving dust case. In particular, the ambipolar resistivity, $\eta_{\mathrm{AD}}$, is very high prior to $n_{\mathrm{H}}=10^{10}\ \mathrm{cm^{-3}}$. As a consequence, magnetic braking is expected to be ineffective. In this case, large size protoplanetary discs should result, which is inconsistent with recent observations. To alleviate this tension, we identified mechanisms that are capable of reducing the ambipolar resistivity during the ensuing protostellar collapse. Among them, electrostatic repulsion and grain-grain erosion feature as the most promising approaches.
Conclusions. The evolution of the magnetic resistivities during the protostellar collapse and consequently the shape of the magnetic field in the early life of the protoplanetary disc strongly depends on the possibility to repopulate the small grains or to prevent their early coagulation. Therefore, it is crucial to better constrain the collision outcomes and the dust grain’s elastic properties, especially the grain’s surface energy based on both theoretical and experimental approaches.
@@
Hydrodynamics; Magnetohydrodynamics (MHD); Turbulence; Protoplanetary discs; Planets and satellites formation.
@@ 1 Introduction
Interstellar dust plays a significant role in many astrophysical environments, in particular, in the context of star formation. Although it represents (on average) only $1\%$ of the diffuse interstellar medium (ISM) mass, the wide spectrum of dust grain sizes affects the optical properties of the medium and, consequently, the heating and cooling processes at work (McKee & Ostriker, 2007). In addition, dust grains are at the heart of many molecule formation processes (including $\mathrm{H_{2}}$ Gould & Salpeter, 1963). They are also the fundamental bricks of planets, which are believed to form rapidly in less than 1 Myr (Testi et al., 2014). The pathway to planetesimals and thereafter planets remains poorly understood. Getting a better description of the dust evolution during the protostellar collapse would allow us to provide more realistic dust initial conditions in protoplanetary discs. Finally, dust grains affect the ionisation degree of the medium and can become the main charge carriers during the protostellar collapse as they allow for electron and ion recombination and electron capture at their surface (Marchand et al., 2016). As such, they control the degree of coupling between the gas and the magnetic field (Nakano et al., 2002) via the ambipolar, Hall, and ohmic magnetic resistivities, and, consequently, the evolution of the angular momentum through magnetic braking (Mouschovias, 1991). Indeed, dust plays a crucial role in shaping the magnetic field during the protostellar collapse and, thus, in influencing the structure of the resulting protoplanetary discs (Wurster et al., 2016, 2018; Hennebelle et al., 2020).
@@
The dust size spectrum is observed or at least modelled as a power-law Mathis-Rumple-Nordsiek (MRN) distribution in the diffuse ISM (Mathis et al., 1977). It was noticed that modifying the population of very small grains in the protostellar collapse simulations radically affects the magnetic resistivities and, consequently, the formation of protoplanetary discs (Zhao et al., 2016; Marchand et al., 2016). In addition, the dust size distribution is expected to evolve as dust grains undergo coagulation and fragmentation processes (Dominik & Tielens, 1997; Ormel et al., 2009). Following such an evolution of the dust distribution is very challenging, but it is essential to correctly describe the non-ideal MHD effects since the coupling between the grains, the gas, and the magnetic field depends directly on the total dust cross section. Indeed, Guillet et al. (2020), Silsbee et al. (2020) and Lebreuilly et al. (2023) showed that a coagulating dust distribution would affect the resistivities during the protostellar collapse. So far, only a handful of studies have accounted for dust coagulation in 2D (Vorobyov & Elbakyan, 2019; Tu et al., 2022) or 3D (Tsukamoto et al., 2021; Bate, 2022; Tsukamoto et al., 2023). Recently, Marchand et al. (2023) performed the first 3D simulations of the protostellar collapse and the early evolution of the protoplanetary disc, while including and self-consistently computing the magnetic field from the magnetic resistivities. The dust coagulation (fragmentation was not included) was modelled at low computational cost with a turbulent kernel (as described in Marchand et al., 2021), exerting an indirect feedback on the gas through the magnetic field, which allows us to better describe the effect of dust growth on non-ideal MHD effects. Similarly, Tsukamoto et al. (2023) used a dust single-size approximation and describe the co-evolution of dust grains and protoplanetary discs. Fragmentation has been introduced by Lebreuilly et al. (2023) and Kawasaki et al. (2022), and some parameters (turbulent intensity and velocity threshold for fragmentation) was investigated in more detail in Kawasaki & Machida (2023). They show that the interplay between coagulation and fragmentation plays a critical role in regard to the evolution of the magnetic resistivities. Indeed, those studies showed that fragmentation could potentially repopulate the small grains at intermediate and high densities, consequently allowing the Ohm and Hall resistivities to rise. The ambipolar resistivity, $\eta_{\mathrm{AD}}$, was reduced to some extent prior to $n_{\mathrm{H}}<10^{10}\mathrm{cm^{-3}}$ and enhanced thereafter. Besides, every study accounting for dust evolution produces magnetic resistivities, which remain remarkably away from the ones obtained with a non-evolving MRN distribution. This is even more striking when including ambipolar diffusion as a source of grain-grain collision, which is very efficient at depleting the smallest grains. For instance, as a consequence of small grain depletion by coagulation, the ambipolar resistivity, $\eta_{\mathrm{AD}}$, exhibits very high values at low and intermediate densities. This would have a profound impact on the properties of the resulting protoplanetary discs. Given that recent observational surveys, including CALYPSO (Maury et al., 2019), have shown that large discs ($>60\ \mathrm{AU}$) are rare, serious doubts are cast on the likeliness of such high ambipolar resistivity values.
@@
To shed light on the impact of dust evolution on the magnetic resistivities during the protostellar collapse, we conducted a thorough investigation of the influence of key parameters and physical effects through single-zone collapse numerical simulations. In particular, we investigated the role of dust grain properties (grain surface energy, intrinsic density, monomer size), collision outcomes (fragments distribution, maximum fragment size, grain-grain erosion, sticking efficiency, electrostatic repulsion), and the collapse properties (dust-to-gas ratio, turbulence intensity, ambipolar drift intensity, initial magnetic field strength, and ionisation rate induced by cosmic rays). However, we focus only on the most influential of the aforementioned parameters and strive to identify mechanisms that would allow us to mitigate the sharp rise in ambipolar resistivity, $\eta_{\mathrm{AD}}$.
@@
In Sects. 2 and 3, we present the dust model and the numerical methods. Then, in Sect. 4 and 5, we share some of our results, which are discussed in Sect. 6. Finally, Sect. 7 provides concluding remarks and additional material can be found in the appendix.
@@ 2.1 Sources of dust relative velocities
Here, we present the different sources of relative velocities that drive dust evolution. In the rest of the paper, the subscripts $i$ and $j$ will refer to the two colliding grains.
@@ 2.1.1 Brownian motion
Dust grains are prone to Brownian motion, which is a thermal process that induces relative velocities as per:
@@ 2.1.2 Ambipolar diffusion
where $t_{\mathrm{s,i}}$ is the stopping time, representing the coupling of a grain to the gas (Epstein, 1924):
@@
where $\rho$ and $c_{\mathrm{s}}\equiv\sqrt{\gamma P/\rho}$ are respectively the mass density of the gas and the sound speed. The number of charges carried by a grain increases slower with respect to grain size than the grain mass does. Therefore, larger grains are associated with lower gyromagnetic frequencies, namely, they are poorly coupled to the magnetic field. Besides, they are weakly coupled to the gas as well, as suggested by the size dependency of the stopping time. Nonetheless, the stopping time and the gyromagnetic timescale both rise with respect to the grain size, but the latter does so more rapidly. This leads in fine to lower Hall factors for larger grains.
As in Spitzer (1941), we define the ambipolar diffusion velocity as:
@@
where $c$ is the speed of light, $\eta_{\mathrm{AD}}$ is the ambipolar resistivity (in unit $\mathrm{s}$). The curl of the magnetic field is approximated as $\frac{|\mathbf{B}|}{\lambda_{\mathrm{J}}}$, where $\lambda_{\mathrm{J}}$ is the Jeans length. Following Guillet et al. (2007, 2020), we have:
@@
where $\delta_{\mathrm{AD}}$ is the amibipolar drift intensity, taken equal to unity if not specified otherwise. We note that the larger the ambipolar resistivity, $\eta_{\mathrm{AD}}$, the stronger the ambipolar drift velocity. Finally, the ambipolar drift velocity between two dust grains is given by:
@@
Ambipolar drift allows for the largest grains to collect the smallest ones. Therefore, ambipolar diffusion is a process very effective at removing the smallest grains, but not at increasing significantly the maximum size of the distribution (Lebreuilly et al., 2023). Importantly, we stress that the ambipolar resistivity, $\eta_{\mathrm{AD}}$, and the ambipolar drift intensity, $\delta_{\mathrm{AD}}$, are two different quantities, which should not be confused henceforth.
@@ 2.1.3 Turbulence
The Stokes number of a grain is defined as $\mathrm{St_{i}}=t_{\mathrm{s,i}}/t_{\mathrm{ff}}$, where the stopping time $t_{\mathrm{s,i}}$ of the largest of the two grains determines the three coupling regimes defined in the framework of Ormel et al. (2009). The differential turbulent induced velocity writes:
@@
where $V_{\mathrm{g}}=\sqrt{\theta}c_{\rm{s}}$ and $\beta_{\mathrm{i,j}}=3.2-(1+x_{i,j})+\frac{2}{1+x_{i,j}}(\frac{1}{2.6}+\frac{x_{i,j}^{3}}{1.6+x_{i,j}})$, with $x_{j,i}$ being the ratio between $\mathrm{St}_{j}$ and $\mathrm{St}_{i}$. The parameter $\theta$ is the turbulence intensity. The usual value taken in the literature in the context of a gravitational collapse is $\theta=\frac{3}{2}$ (Guillet et al., 2020).
When the stopping time of the largest grain is comprised between the dissipation timescale and the injection timescale (i.e. turnover time of the largest eddies), the contribution from class II eddies acts as a random kick in the grain motion, allowing for grains of similar size to efficiently collide (see Ormel & Cuzzi, 2007, for a detailed definition of the class I, II and III eddies). This is the intermediate regime, which is the dominant regime in the context of the protostellar collapse (Marchand et al., 2021).
We note that the turbulent source of relative velocities is commonly found to be the most efficient at growing large grains and significantly increasing the maximum size of the dust distribution (Silsbee et al., 2020; Guillet et al., 2020; Lebreuilly et al., 2023).
@@ 2.2 Dust evolution
The different sources of relative velocity presented here lead to grain collisions that may result in various outcomes. In this work, we only account for grain coagulation (perfect and imperfect sticking) and fragmentation, and we also investigate the influence of grain-grain erosion with a simplistic approach. Many other mechanisms have been identified with laboratory experiments, such as gas-grain erosion, abrasion, compaction and bouncing, mass transfer, and rotational disruption (see Wurm & Teiser, 2021, for more details). The grains are considered perfectly spherical, with a fixed and unique intrinsic density.
@@ 2.2.1 Coagulation
The equation of mass conservation for the grains is expressed as:
@@
The subscript $k$ designates one specific grain among $\mathcal{N}$ species. Here, the term species relates to a grain of given size. The source term $S_{k,\mathrm{growth}}$ accounts for coagulation and fragmentation events and is computed according to the Smoluchowski equation (Smoluchowski, 1916):
@@
where $\Delta v_{i,j}$ is the differential velocity between grains $i$ and $j$, of size $s_{i}$ and $s_{j}$, defined as the quadratic sum of the sources of differential velocity detailed in Sect. 2.1. The $\sqrt{\frac{8}{3\pi}}$ pre-factor comes from considering that grains relative velocities along the three x, y, and z axis are Gaussian variables of variance $\Delta v_{i,j}^{2}/3$ (Guillet et al., 2020; Marchand et al., 2021). Here, $E_{\mathrm{kinetic}}=\frac{1}{2}\mu_{i,j}\Delta v_{i,j}^{2}$ is the kinetic energy of the colliding grains. $\mu_{i,j}\equiv m_{i}m_{j}/\left(m_{i}+m_{j}\right)$ is the reduced mass of the two grains. Also, $E_{\mathrm{Coulomb}}=q_{\mathrm{i}}q_{\mathrm{j}}/4\pi\epsilon_{\mathrm{vacuum}}\left(s_{\mathrm{i}}+s_{\mathrm{j}}\right)$ is the Coulomb electrostatic energy and $\epsilon_{\mathrm{vacuum}}$ the vacuum dielectric permittivity. The first term in the Kernel accounts for the cross section modification induced by the electrostatic interaction between the two colliding grains (Akimkin et al., 2023). It is included only in the simulations for which electrostatic repulsion is explicitly investigated.